Solving:
-2 x-3y = 14
5 x-5y = 15

Using Substitution Method.

Answer by siddhant(1)

-2x=14+3y
x=-(14+3y)\2
therefore, substitute the value of x in equation 5x-5y=15
5(-14-3y)\2-5y=15
[(-70-15y)-10y]/2=15
(-25y-70)=30
y=-4
x=-1